A small, unshaded lamp of brightness 400 candelas, which may be considered a point source of light, is placed 4 meters above the floor of a big hall. Find: 


A. the total luminous flux emitted by the source, and

B. the intensity of illumination on the floor directly below the lamp.



1
Expert's answer
2021-04-12T06:58:36-0400

a)


"\Phi=4\pi(400)=5027\ Lumen"

b)

"E=\frac{400}{4^2}=25\frac{cd}{m^2}"